# Are foundation models for computer vision good conformal predictors?

### v2

@@ G Few-Shot Adaptation
Table 10: Set size and CovGap for CLIP (ViT-B backbone) on 11 datasets across several few-shot adaptation approaches. Training for 4 shots.
@@
Aircraft ZS 18.7 19.75 20.74 0.138 0.139 0.127
ZSLP 8.14 9.56 9.43 0.079 0.069 0.082
CLAP [53] 8.20 9.59 9.43 0.079 0.070 0.083
CoOp [70] 7.19 8.09 8.88 0.080 0.070 0.084
KgCoOp [17] 7.53 8.72 9.13 0.080 0.075 0.084
Caltech-101 ZS 0.94 1.95 1.20 0.136 0.092 0.092
ZSLP 0.91 1.12 1.02 0.151 0.083 0.082
CLAP [53] 0.91 1.22 1.06 0.146 0.083 0.085
CoOp [70] 0.91 1.06 1.02 0.148 0.084 0.085
KgCoOp [17] 0.91 1.19 1.06 0.150 0.082 0.083
DTD ZS 11.78 13.3 13.66 0.122 0.117 0.121
ZSLP 2.91 4.90 3.34 0.081 0.071 0.073
CLAP [53] 3.03 5.12 3.48 0.080 0.069 0.079
CoOp [70] 3.37 4.50 3.94 0.084 0.076 0.081
KgCoOp [17] 3.53 4.81 4.04 0.082 0.073 0.082
EuroSAT ZS 4.55 4.78 5.06 0.103 0.101 0.102
ZSLP 1.46 2.04 1.84 0.053 0.045 0.038
CLAP [53] 1.53 2.09 1.90 0.051 0.044 0.038
CoOp [70] 1.23 1.59 1.50 0.041 0.026 0.031
KgCoOp [17] 1.28 1.73 1.61 0.045 0.032 0.032
Flowers-102 ZS 9.26 9.52 15.15 0.180 0.175 0.179
ZSLP 1.03 2.19 1.58 0.132 0.087 0.091
CLAP [53] 1.09 2.30 1.65 0.135 0.089 0.094
CoOp [70] 0.93 1.30 1.17 0.094 0.078 0.078
KgCoOp [17] 0.93 1.44 1.24 0.102 0.079 0.078
Food-101 ZS 1.14 1.89 1.45 0.054 0.026 0.031
ZSLP 1.08 1.77 1.39 0.044 0.023 0.025
CLAP [53] 1.07 1.79 1.39 0.045 0.023 0.024
CoOp [70] 1.11 1.64 1.35 0.050 0.025 0.03
KgCoOp [17] 1.08 1.75 1.38 0.048 0.022 0.026
ImageNet ZS 2.81 10.07 3.23 0.086 0.069 0.084
ZSLP 2.17 6.60 2.44 0.076 0.060 0.074
CLAP [53] 2.15 6.85 2.45 0.077 0.060 0.074
CoOp [70] 2.39 5.27 2.73 0.078 0.064 0.077
KgCoOp [17] 2.30 5.95 2.64 0.078 0.062 0.077
Oxford Pets ZS 1.05 1.43 1.33 0.109 0.067 0.072
ZSLP 0.94 1.24 1.17 0.081 0.040 0.037
CLAP [53] 0.95 1.27 1.19 0.089 0.039 0.038
CoOp [70] 0.94 1.16 1.13 0.083 0.040 0.041
KgCoOp [17] 0.94 1.20 1.15 0.083 0.038 0.038
Stanford Cars ZS 2.24 3.09 2.49 0.109 0.08 0.102
ZSLP 1.27 1.98 1.66 0.082 0.058 0.057
CLAP [53] 1.32 2.04 1.7 0.080 0.057 0.057
CoOp [70] 1.37 1.94 1.66 0.078 0.057 0.059
KgCoOp [17] 1.39 2.04 1.73 0.080 0.059 0.061
SUN397 ZS 2.83 5.80 3.21 0.093 0.073 0.093
ZSLP 1.72 3.49 2.09 0.075 0.055 0.063
CLAP [53] 1.76 3.55 2.12 0.075 0.056 0.064
CoOp [70] 1.87 2.73 2.04 0.073 0.060 0.072
KgCoOp [17] 1.85 2.91 2.07 0.076 0.059 0.070
UCF101 ZS 2.89 5.24 3.83 0.124 0.096 0.124
ZSLP 1.36 2.85 1.82 0.106 0.068 0.077
CLAP [53] 1.41 2.94 1.87 0.108 0.067 0.079
CoOp [70] 1.51 2.28 1.76 0.108 0.068 0.095
KgCoOp [17] 1.41 2.47 1.77 0.106 0.067 0.081
Table 11: Set size and CovGap for CLIP (ViT-B backbone) on 11 datasets across several few-shot adaptation approaches. Training for 16 shots.

### v3

@@ Appendix F Few-Shot Adaptation
Table 13: Set size and CovGap for CLIP (ViT-B) on 11 datasets across several few-shot adaptation approaches. Training for 16 shots.
@@
Aircraft ZS 18.70 $\pm$ 0.06 19.75 $\pm$ 0.06 20.74 $\pm$ 0.05 0.138 $\pm$ 0.000 0.139 $\pm$ 0.000 0.127 $\pm$ 0.000
ZSLP 8.14 $\pm$ 0.03 9.56 $\pm$ 0.03 9.43 $\pm$ 0.04 0.079 $\pm$ 0.001 0.069 $\pm$ 0.000 0.082 $\pm$ 0.001
CoOp (Zhou et al., 2022b) 7.19 $\pm$ 0.02 8.09 $\pm$ 0.02 8.88 $\pm$ 0.04 0.080 $\pm$ 0.001 0.070 $\pm$ 0.000 0.084 $\pm$ 0.001
KgCoOp (Hantao Yao, 2023) 7.53 $\pm$ 0.02 8.72 $\pm$ 0.03 9.13 $\pm$ 0.04 0.080 $\pm$ 0.001 0.075 $\pm$ 0.000 0.084 $\pm$ 0.001
CLAP (Silva-Rodríguez et al., 2024) 8.20 $\pm$ 0.03 9.59 $\pm$ 0.03 9.43 $\pm$ 0.04 0.079 $\pm$ 0.001 0.070 $\pm$ 0.000 0.083 $\pm$ 0.001
Caltech-101 ZS 0.94 $\pm$ 0.00 1.95 $\pm$ 0.01 1.20 $\pm$ 0.00 0.136 $\pm$ 0.001 0.092 $\pm$ 0.001 0.092 $\pm$ 0.001
ZSLP 0.91 $\pm$ 0.00 1.12 $\pm$ 0.00 1.02 $\pm$ 0.00 0.151 $\pm$ 0.001 0.083 $\pm$ 0.001 0.082 $\pm$ 0.001
CoOp (Zhou et al., 2022b) 0.91 $\pm$ 0.00 1.06 $\pm$ 0.00 1.02 $\pm$ 0.00 0.148 $\pm$ 0.001 0.084 $\pm$ 0.001 0.085 $\pm$ 0.001
KgCoOp (Hantao Yao, 2023) 0.91 $\pm$ 0.00 1.19 $\pm$ 0.00 1.06 $\pm$ 0.00 0.150 $\pm$ 0.001 0.082 $\pm$ 0.001 0.083 $\pm$ 0.001
CLAP (Silva-Rodríguez et al., 2024) 0.91 $\pm$ 0.00 1.22 $\pm$ 0.01 1.06 $\pm$ 0.00 0.146 $\pm$ 0.001 0.083 $\pm$ 0.001 0.085 $\pm$ 0.001
DTD ZS 11.78 $\pm$ 0.06 13.30 $\pm$ 0.04 13.66 $\pm$ 0.06 0.122 $\pm$ 0.001 0.117 $\pm$ 0.001 0.121 $\pm$ 0.001
ZSLP 2.91 $\pm$ 0.01 4.90 $\pm$ 0.03 3.34 $\pm$ 0.02 0.081 $\pm$ 0.001 0.071 $\pm$ 0.001 0.073 $\pm$ 0.001
CoOp (Zhou et al., 2022b) 3.37 $\pm$ 0.02 4.50 $\pm$ 0.03 3.94 $\pm$ 0.04 0.084 $\pm$ 0.001 0.076 $\pm$ 0.001 0.081 $\pm$ 0.001
KgCoOp (Hantao Yao, 2023) 3.53 $\pm$ 0.02 4.81 $\pm$ 0.03 4.04 $\pm$ 0.04 0.082 $\pm$ 0.001 0.073 $\pm$ 0.001 0.082 $\pm$ 0.001
CLAP (Silva-Rodríguez et al., 2024) 3.03 $\pm$ 0.02 5.12 $\pm$ 0.03 3.48 $\pm$ 0.02 0.080 $\pm$ 0.001 0.069 $\pm$ 0.001 0.079 $\pm$ 0.001
EuroSAT ZS 4.55 $\pm$ 0.01 4.78 $\pm$ 0.00 5.06 $\pm$ 0.01 0.103 $\pm$ 0.001 0.101 $\pm$ 0.001 0.102 $\pm$ 0.000
ZSLP 1.46 $\pm$ 0.00 2.04 $\pm$ 0.00 1.84 $\pm$ 0.00 0.053 $\pm$ 0.000 0.045 $\pm$ 0.000 0.038 $\pm$ 0.000
CoOp (Zhou et al., 2022b) 1.23 $\pm$ 0.00 1.59 $\pm$ 0.00 1.50 $\pm$ 0.00 0.041 $\pm$ 0.000 0.026 $\pm$ 0.000 0.031 $\pm$ 0.000
KgCoOp (Hantao Yao, 2023) 1.28 $\pm$ 0.00 1.73 $\pm$ 0.00 1.61 $\pm$ 0.00 0.045 $\pm$ 0.000 0.032 $\pm$ 0.000 0.032 $\pm$ 0.000
CLAP (Silva-Rodríguez et al., 2024) 1.53 $\pm$ 0.00 2.09 $\pm$ 0.00 1.90 $\pm$ 0.00 0.051 $\pm$ 0.000 0.044 $\pm$ 0.000 0.038 $\pm$ 0.000
Flowers-102 ZS 9.26 $\pm$ 0.05 9.52 $\pm$ 0.06 15.15 $\pm$ 0.08 0.180 $\pm$ 0.000 0.175 $\pm$ 0.000 0.179 $\pm$ 0.000
ZSLP 1.03 $\pm$ 0.00 2.19 $\pm$ 0.01 1.58 $\pm$ 0.00 0.132 $\pm$ 0.001 0.087 $\pm$ 0.001 0.091 $\pm$ 0.001
CoOp (Zhou et al., 2022b) 0.93 $\pm$ 0.00 1.30 $\pm$ 0.00 1.17 $\pm$ 0.00 0.094 $\pm$ 0.001 0.078 $\pm$ 0.001 0.078 $\pm$ 0.001
KgCoOp (Hantao Yao, 2023) 0.93 $\pm$ 0.00 1.44 $\pm$ 0.01 1.24 $\pm$ 0.00 0.102 $\pm$ 0.001 0.079 $\pm$ 0.000 0.078 $\pm$ 0.001
CLAP (Silva-Rodríguez et al., 2024) 1.09 $\pm$ 0.00 2.30 $\pm$ 0.01 1.65 $\pm$ 0.00 0.135 $\pm$ 0.001 0.089 $\pm$ 0.001 0.094 $\pm$ 0.001
Food-101 ZS 1.14 $\pm$ 0.00 1.89 $\pm$ 0.00 1.45 $\pm$ 0.00 0.054 $\pm$ 0.000 0.026 $\pm$ 0.000 0.031 $\pm$ 0.000
ZSLP 1.08 $\pm$ 0.00 1.77 $\pm$ 0.00 1.39 $\pm$ 0.00 0.044 $\pm$ 0.000 0.023 $\pm$ 0.000 0.025 $\pm$ 0.000
CoOp (Zhou et al., 2022b) 1.11 $\pm$ 0.00 1.64 $\pm$ 0.00 1.35 $\pm$ 0.00 0.050 $\pm$ 0.000 0.025 $\pm$ 0.000 0.030 $\pm$ 0.000
KgCoOp (Hantao Yao, 2023) 1.08 $\pm$ 0.00 1.75 $\pm$ 0.00 1.38 $\pm$ 0.00 0.048 $\pm$ 0.000 0.022 $\pm$ 0.000 0.026 $\pm$ 0.000
CLAP (Silva-Rodríguez et al., 2024) 1.07 $\pm$ 0.00 1.79 $\pm$ 0.00 1.39 $\pm$ 0.00 0.045 $\pm$ 0.000 0.023 $\pm$ 0.000 0.024 $\pm$ 0.000
ImageNet ZS 2.81 $\pm$ 0.00 10.07 $\pm$ 0.02 3.23 $\pm$ 0.00 0.086 $\pm$ 0.000 0.069 $\pm$ 0.000 0.084 $\pm$ 0.000
ZSLP 2.17 $\pm$ 0.00 6.60 $\pm$ 0.01 2.44 $\pm$ 0.00 0.076 $\pm$ 0.000 0.060 $\pm$ 0.000 0.074 $\pm$ 0.000
CoOp (Zhou et al., 2022b) 2.39 $\pm$ 0.00 5.27 $\pm$ 0.01 2.73 $\pm$ 0.00 0.078 $\pm$ 0.000 0.064 $\pm$ 0.000 0.077 $\pm$ 0.000
KgCoOp (Hantao Yao, 2023) 2.30 $\pm$ 0.00 5.95 $\pm$ 0.01 2.64 $\pm$ 0.00 0.078 $\pm$ 0.000 0.062 $\pm$ 0.000 0.077 $\pm$ 0.000
CLAP (Silva-Rodríguez et al., 2024) 2.15 $\pm$ 0.00 6.85 $\pm$ 0.01 2.45 $\pm$ 0.00 0.077 $\pm$ 0.000 0.060 $\pm$ 0.000 0.074 $\pm$ 0.000
Oxford Pets ZS 1.05 $\pm$ 0.00 1.43 $\pm$ 0.00 1.33 $\pm$ 0.00 0.109 $\pm$ 0.000 0.067 $\pm$ 0.000 0.072 $\pm$ 0.000
ZSLP 0.94 $\pm$ 0.00 1.24 $\pm$ 0.00 1.17 $\pm$ 0.00 0.081 $\pm$ 0.001 0.040 $\pm$ 0.000 0.037 $\pm$ 0.000
CoOp (Zhou et al., 2022b) 0.94 $\pm$ 0.00 1.16 $\pm$ 0.00 1.13 $\pm$ 0.00 0.083 $\pm$ 0.000 0.040 $\pm$ 0.000 0.041 $\pm$ 0.000
KgCoOp (Hantao Yao, 2023) 0.94 $\pm$ 0.00 1.20 $\pm$ 0.00 1.15 $\pm$ 0.00 0.083 $\pm$ 0.001 0.038 $\pm$ 0.001 0.038 $\pm$ 0.000
CLAP (Silva-Rodríguez et al., 2024) 0.95 $\pm$ 0.00 1.27 $\pm$ 0.00 1.19 $\pm$ 0.00 0.089 $\pm$ 0.001 0.039 $\pm$ 0.000 0.038 $\pm$ 0.001
Cars ZS 2.24 $\pm$ 0.00 3.09 $\pm$ 0.01 2.49 $\pm$ 0.00 0.109 $\pm$ 0.000 0.080 $\pm$ 0.000 0.102 $\pm$ 0.000
ZSLP 1.27 $\pm$ 0.00 1.98 $\pm$ 0.00 1.66 $\pm$ 0.00 0.082 $\pm$ 0.000 0.058 $\pm$ 0.000 0.057 $\pm$ 0.000
CoOp (Zhou et al., 2022b) 1.37 $\pm$ 0.00 1.94 $\pm$ 0.00 1.66 $\pm$ 0.00 0.078 $\pm$ 0.000 0.057 $\pm$ 0.000 0.059 $\pm$ 0.000
KgCoOp (Hantao Yao, 2023) 1.39 $\pm$ 0.00 2.04 $\pm$ 0.00 1.73 $\pm$ 0.00 0.080 $\pm$ 0.000 0.059 $\pm$ 0.000 0.061 $\pm$ 0.000
CLAP (Silva-Rodríguez et al., 2024) 1.32 $\pm$ 0.00 2.04 $\pm$ 0.00 1.70 $\pm$ 0.00 0.080 $\pm$ 0.000 0.057 $\pm$ 0.000 0.057 $\pm$ 0.000
SUN397 ZS 2.83 $\pm$ 0.01 5.80 $\pm$ 0.01 3.21 $\pm$ 0.01 0.093 $\pm$ 0.000 0.073 $\pm$ 0.000 0.093 $\pm$ 0.000
ZSLP 1.72 $\pm$ 0.00 3.49 $\pm$ 0.01 2.09 $\pm$ 0.00 0.075 $\pm$ 0.000 0.055 $\pm$ 0.000 0.063 $\pm$ 0.000
CoOp (Zhou et al., 2022b) 1.87 $\pm$ 0.00 2.73 $\pm$ 0.00 2.04 $\pm$ 0.00 0.073 $\pm$ 0.000 0.060 $\pm$ 0.000 0.072 $\pm$ 0.000
KgCoOp (Hantao Yao, 2023) 1.85 $\pm$ 0.00 2.91 $\pm$ 0.01 2.07 $\pm$ 0.00 0.076 $\pm$ 0.000 0.059 $\pm$ 0.000 0.070 $\pm$ 0.000
CLAP (Silva-Rodríguez et al., 2024) 1.76 $\pm$ 0.00 3.55 $\pm$ 0.01 2.12 $\pm$ 0.00 0.075 $\pm$ 0.000 0.056 $\pm$ 0.000 0.064 $\pm$ 0.000
UCF101 ZS 2.89 $\pm$ 0.01 5.24 $\pm$ 0.02 3.83 $\pm$ 0.01 0.124 $\pm$ 0.000 0.096 $\pm$ 0.000 0.124 $\pm$ 0.001
ZSLP 1.36 $\pm$ 0.00 2.85 $\pm$ 0.01 1.82 $\pm$ 0.00 0.106 $\pm$ 0.001 0.068 $\pm$ 0.000 0.077 $\pm$ 0.000
CoOp (Zhou et al., 2022b) 1.51 $\pm$ 0.00 2.28 $\pm$ 0.01 1.76 $\pm$ 0.00 0.108 $\pm$ 0.001 0.068 $\pm$ 0.000 0.095 $\pm$ 0.000
KgCoOp (Hantao Yao, 2023) 1.41 $\pm$ 0.00 2.47 $\pm$ 0.01 1.77 $\pm$ 0.00 0.106 $\pm$ 0.001 0.067 $\pm$ 0.000 0.081 $\pm$ 0.000
CLAP (Silva-Rodríguez et al., 2024) 1.41 $\pm$ 0.00 2.94 $\pm$ 0.01 1.87 $\pm$ 0.00 0.108 $\pm$ 0.001 0.067 $\pm$ 0.001 0.079 $\pm$ 0.000